OpenStudy (cwrw238):

What's the next number in this series: 0, 1 , 5, 23, 119, _

OpenStudy (cwrw238):

clue: this involves subtraction

OpenStudy (anonymous):

really? how?

OpenStudy (cwrw238):

subtract 1 from each term

OpenStudy (cwrw238):

got it ganeshie8

ganeshie8 (ganeshie8):

deleted ;p

OpenStudy (jhannybean):

xD I had the factorial but i got confused after!!!!! noooo!!!! lol.

OpenStudy (cwrw238):

next term is 6! - 1

OpenStudy (cwrw238):

not easy, this one

OpenStudy (anonymous):

what? haha..

OpenStudy (kropot72):

\[a _{n}=n!-1\]

OpenStudy (jhannybean):

yaaaay

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

ahh.. :)) i got it. :)) ^_____^ thanks.

OpenStudy (jhannybean):

Hmm..it's the opposite of the other kind of..? the 110 is throwing me off -_-

OpenStudy (cwrw238):

clue: different base

OpenStudy (anonymous):

ano daw? -.-

OpenStudy (cwrw238):

ooopppss sorry i/ve made a mistake correct is 110, 20 , 12, 11, 10, _

OpenStudy (jhannybean):

>:|

OpenStudy (cwrw238):

each term can be converted to the same number

OpenStudy (cwrw238):

right

OpenStudy (jhannybean):

ohh imkind of seeing it... 110 = 11, 10 sort of?

OpenStudy (cwrw238):

no 110 = 6 20 = 6 12 = 6 11 = 6

OpenStudy (jhannybean):

Wait are these 6's...i was just looking at a pattern i noticed after staring at these numbers for a good couple mins xD

OpenStudy (cwrw238):

110 is 6 in binary system 20 is 6 in the number system based on 3 ( 2*3 + 0 = 6) 12 is 6 in .......................................... 4 ( 1*4 + 2 = 6) follow?

ganeshie8 (ganeshie8):

next is 6 in base7.. .this is real nice :)

OpenStudy (cwrw238):

10 = 6 in number susten based on 6 ( 1*6 + 0) finally 6 = 6 system based on 7

OpenStudy (jhannybean):

haha im reading on this and comparing it to the problem he posted.Never learned it yet but it looks interesting.

OpenStudy (cwrw238):

we humans use base10 system maybe because we have ten fingers

hartnn (hartnn):

thats one reason to use base 10, yes. Any number can be converted to base 10 using this :   and so on...